find the height of the projectile at highest point ,if it is projected with speed 15m/s ,in the direction 45​*(45 degrees) above the horizontal ?

Grade:11

1 Answers

Shashank Raut
7 years ago
Use H=u²sin²x/2g where x is angle of projectionH=15*15*0.5/2*10 ( 0.5 is from sin²45)H=45/8=5.625 metre